OpenStudy (anonymous):

absoulute value equations |7+3a|=11-a

OpenStudy (anonymous):

Split this into 2 equations. Since what's in the absolute value sticks could be either positive or negative, you have to take that into account. So... 7+3a=11-a or -(7+3a)=11-a Then solve for a.